OpenStudy (anonymous):

I will give medal to who ever answers these!!! What are the missing y-values in the table below? y = x - 2 x y 3 ? 6 ? 8 ? 10 ?

OpenStudy (mathstudent55):

Substitute the x in the equation with each given x value, and calculate y.

OpenStudy (anonymous):

What is the missing numbers for 3-5-8-10 and what is the 1st x and y? Help please

OpenStudy (anonymous):

?

OpenStudy (mathstudent55):

For x = 3, you do this: y = x - 2 Substitute x with 3: y = 3 - 2 y = 1 That means for x = 3, y = 1. Now do the same for the other values of x.

OpenStudy (anonymous):

Oh ok thank you

OpenStudy (mathstudent55):

You're welcome.